OpenStudy (anonymous):

Please help and explain!!!!!! I will fan and give medals! If f(x)= 3/(x+1), which equations represents the inverse of function f(x)? f^-1(x)= -3/(x+1) f^-1(x)= 3/(x-1) f^-1(x)= (x+1)/3 f^-1(x)= 3x/(1+x)

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

you are looking for the inverse function

OpenStudy (anonymous):

\[f(x)=\frac{3}{x+1}\] and you want \(f^{-1}(x)\) the inverse start by setting \[y=\frac{3}{x+1}\] then switch \(x\) and \(y\) to get \[x=\frac{3}{y+1}\] and solve the equation for \(y\)

OpenStudy (anonymous):

this takes a couple steps, do you know how to do it?

OpenStudy (anonymous):

I think so. Is the correct answer 

OpenStudy (anonymous):

i don't think so lets solve it

OpenStudy (anonymous):

damn typo \[x=\frac{3}{y+1}\] \[x(y+1)=3\] \[xy+x=3\] \[xy=3-x\] \[y=\frac{3-x}{x}\]

OpenStudy (anonymous):

so answer is \[f^{-1}(x)=\frac{3-x}{x}\]

OpenStudy (anonymous):

Thank you so much! Now I know for sure how to solve problems like this. You helped tremendously!

OpenStudy (anonymous):

yw i notice however that this is not one of your answer choices i am sure that it is right, but i am confused as to why this doesn't appear in your list